# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v3

@@ III.3 Total 3D Hall conductivity
Figure 5: Total 3D Hall conductivity [Eq. (8)] as a function of the interlayer hopping strength $\eta_{z}$. (a) Full range: $\eta_{z}/t_{0}\in[0,2]$; (b) Enlarged view of the moderate-coupling regime: $\eta_{z}/t_{0}\in[0,0.2]$. Here, $L_{z}\!=\!\lambda_{z}\!=\!61a$, and the other parameters are the same as those used in Fig. 3.
@@
To investigate the effects of interlayer coupling on the total 3D Hall conductivity, we compute it as a function of the interlayer hopping strength $\eta_{z}$, as shown in Fig. 5(a), based on Eq. (8). The results indicate that for $\eta_{z}/t_{0}>0.05$, the total 3D Hall conductivity decreases monotonically as $\eta_{z}$ increases, eventually approaching zero when $\eta_{z}\sim 2t_{0}$. This behavior indicates that strong interlayer coupling can effectively suppress the total 3D Hall conductivity.
@@
In particular, a closer look at the range $\eta_{z}/t_{0}\in[0,0.2]$ [Fig. 5(b)] reveals an approximate plateau within $\eta_{z}/t_{0}\in[0,0.05]$. This suggests that the total 3D Hall conductivity is robust against weak interlayer coupling.
IV Floquet Hamiltonian
@@
The optical field propagating along the $z$-direction in the material is given by ${\bf E}(t)\!=\!\partial{\bf A}(t)/\partial t\!=\!E_{0}(\cos(\omega t),\cos(\omega t\!+\!\varphi),0)$, where $E_{0}$ is the amplitude, $\omega$ is the angular frequency, and $\varphi$ represents the phase. For linear polarization, $\varphi\!=\!0$, while $\varphi\!=\!\mp\pi/2$ corresponds to left- or right-handed circular polarization. The vector potential is periodic with period $T\!=\!2\pi/\omega$, and it is given by ${\bf A}(t)\!=\!{\bf A}(t\!+\!T)\!=\!\omega^{-1}E_{0}(\sin(\omega t),\sin(\omega t\!+\!\varphi),0)$. In the off-resonant regime, where the central Floquet band is well-separated from other replicas, the high-frequency expansion applies. We choose the optical frequency as $\hbar\omega\!=\!10$ eV, which is much larger than the material’s bandwidth Zhan et al. (2024); Qin et al. (2023, 2022b, 2022a, 2024); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018).
@@
[…]
where $\mathcal{T}$ is the time-ordering operator. In the high-frequency regime, a closed-form solution exists via the high-frequency expansion Magnus (1954); Blanes et al. (2009); Lee et al. (2018); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, d); Du et al. (2022); Wang et al. (2023):
@@
[…]
IV.1 Floquet Weyl-like model
@@
${\cal J}_{n}(A_{0}a)$ is the $n$th Bessel function of the first kind Temme (1996), and $A_{0}\!=\!eE_{0}/(\hbar\omega)$. The detailed derivations leading to Eq. (24) are provided in Subsection SV A in the Supplemental Material Sup . For $\varphi\!=\!0$ (linearly polarized laser), the Weyl-like nodes are determined by solving the equation $m_{z}\!+\!V_{z}\cos(Q_{z}z)\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\!=\!0$, subject to the condition $-1\leqslant\cos(Q_{z}z)\!=\!\{m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\}/V_{z}\leqslant 1$. This yields the Weyl-like nodes at
@@
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/V_{z}\leqslant 1$. By substituting the model parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $a\!=\!1$ nm into the condition $|\tilde{m}_{z}(A_{0})/V_{z}|<1$, we have $A_{0}<1.034$ nm-1. From Eq. (28), it is clear that the positions of the Weyl-like nodes can be tuned by adjusting the intensity of the optical field.
Furthermore, substituting Eq. (28) into Eq. (9), we can obtain the total 3D Hall conductivity for the Floquet Weyl-like model (24) as
@@
Additionally, Section SVI of the Supplemental Material Sup presents the derivations for the Weyl-like model under PBCs along the $x$-direction and OBCs along the $y$-direction.
IV.2 Floquet Weyl semimetal
The Floquet Hamiltonian for the Weyl semimetal, derived from Eq. (23), takes the form:
@@ IV.2 Floquet Weyl semimetal
A detailed derivation of Eq. (30) is provided in Subsection SV B in the Supplemental Material Sup . For linearly polarized light ($\varphi\!=\!0$), the Weyl nodes are determined by solving the equation $-m_{z}\!+\!t_{z}\cos(k_{z}a)\!-\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\!=\!0$, subject to the condition $-1\leqslant\cos(k_{z}a)\!=\!\{m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\}/t_{z}\leqslant 1$. This condition gives the Weyl node positions:
@@
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/t_{z}\leqslant 1$. Substituting typical model parameters, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $a\!=\!1$ nm, into the condition $|\tilde{m}_{z}(A_{0})/t_{z}|<1$, we obtain the constraint $A_{0}<1.034$ nm-1. From Eq. (34), we can conclude that the positions of the Weyl nodes can be tuned by adjusting the intensity of the optical field.
Furthermore, substituting Eq. (34) into Eq. (15), we can obtain the total 3D Hall conductivity for the Floquet Weyl semimetal (30) as
@@
The detailed derivation of the Floquet Hamiltonian matrix for the Weyl semimetal, under OBCs along the $z$-direction and PBCs along the $x$- and $y$-directions, is provided in Subsection SVII A of the Supplemental Material Sup . Similarly, Subsection SVII B of the Supplemental Material Sup presents the derivation for the case of PBCs along the $x$- and $z$-directions and OBCs along the $y$-direction.
IV.3 Validity

### v4

@@ III.3 Total 3D Hall conductivity
Figure 7: Total 3D Hall conductivity [Eq. (8)] as a function of the interlayer hopping strength $\eta_{z}$. (a) Full range: $\eta_{z}/t_{0}\in[0,2]$. (b) Enlarged view of the moderate-coupling regime: $\eta_{z}/t_{0}\in[0,0.2]$. Here, $L_{z}\!=\!\lambda_{z}\!=\!61a$, and the other parameters are the same as those used in Fig. 5.
@@
To investigate the effects of interlayer coupling on the total 3D Hall conductivity, we compute it as a function of the interlayer hopping strength $\eta_{z}$, as shown in Fig. 7(a), based on Eq. (8). The results indicate that, for $\eta_{z}/t_{0}>0.05$, the total 3D Hall conductivity decreases monotonically as $\eta_{z}$ increases, eventually approaching zero when $\eta_{z}\sim 2t_{0}$. This behavior indicates that strong interlayer coupling can effectively suppress the total 3D Hall conductivity.
@@
A closer look at the range $\eta_{z}/t_{0}\in[0,0.2]$ [Fig. 7(b)] reveals an approximate plateau within $\eta_{z}/t_{0}\in[0,0.05]$. This suggests that the total 3D Hall conductivity is robust against weak interlayer coupling.
IV Disorder
To examine the effect of a generic disorder potential on the Weyl-like points and the associated topological invariant, we introduce an Anderson-type disorder Li et al. (2009); Groth et al. (2009); Stützer et al. (2018); Meier et al. (2018); Li et al. (2020); Qin et al. (2023b); Chen et al. (2024b, 2017a, 2017b, 2018a, 2018b); Yi et al. (2024); Chen and Zhou (2025)
$H_{d}\!=\!\sum_{j_{x},j_{y},j_{z}}W(j_{x},j_{y},j_{z})\sigma_{0}C_{j_{x},j_{y},j_{z}}^{\dagger}C_{j_{x},j_{y},j_{z}},$ (22)
where $C_{j_{x},j_{y},j_{z}}^{{\dagger}}$ ($C_{j_{x},j_{y},j_{z}}$) denotes the electron creation (annihilation) operator at site $(j_{x},j_{y},j_{z})$, the summation runs over all lattice sites, and $\sigma_{0}$ is a $2\times 2$ identity matrix. The onsite disorder potential $W(j_{x},j_{y},j_{z})$ is a random variable uniformly distributed in the energy interval [U,U][-U,U], where $U$ measures the disorder strength.
Based on the model Hamiltonian in Eq. (1), we emphasize that the system is decoupled along the $z$ direction and can therefore be decomposed into multiple 2D models for computation.
To incorporate the Anderson-type disorder in Eq. (22) into our model Hamiltonian, we rewrite the Hamiltonian in real space as follows. By performing 2D Fourier transforms, we obtain the real-space disordered Weyl-like model Hamiltonian with PBCs in the $x$ and $y$ directions. In the basis
$(C_{1,1,j_{z}},C_{1,2,j_{z}},\cdots,C_{1,L_{y},j_{z}},C_{2,1,j_{z}},C_{2,2,j_{z}},\cdots,C_{2,L_{y},j_{z}},$
$\cdots,C_{L_{x},L_{y},j_{z}})^{T},$
the Hamiltonian reads
${\cal H}_{yx}\!=\!\begin{pmatrix}h_{1}&T_{1}&0&\cdots&T_{1}^{\dagger}\\ T_{1}^{\dagger}&h_{1}&T_{1}&\cdots&0\\ 0&T_{1}^{\dagger}&h_{1}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&T_{1}\\ T_{1}&\cdots&0&T_{1}^{\dagger}&h_{1}\end{pmatrix}_{(2L_{y}L_{x})\times(2L_{y}L_{x})},$ (23)
where
$h_{1}\!=\!\begin{pmatrix}M(z)&T_{y}&0&\cdots&T_{y}^{\dagger}\\ T_{y}^{\dagger}&M(z)&T_{y}&\cdots&0\\ 0&T_{y}^{\dagger}&M(z)&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&T_{y}\\ T_{y}&\cdots&0&T_{y}^{\dagger}&M(z)\end{pmatrix}_{2L_{y}\times 2L_{y}},$ (24)
$T_{1}\!=\!\begin{pmatrix}T_{x}&0&0&\cdots&0\\ 0&T_{x}&0&\cdots&0\\ 0&0&T_{x}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&0\\ 0&\cdots&0&0&T_{x}\end{pmatrix}_{2L_{y}\times 2L_{y}}.$ (25)
The terms $M(z)$, $T_{x}$, and $T_{y}$ are given by
$M(z)\!=\![m_{z}\!+\!V(z)\!+\!2m_{0}]\sigma_{z}\!+\!W(j_{x},j_{y},j_{z})\sigma_{0},$ (26)
$T_{x}\!=\!-\frac{m_{0}}{2}\sigma_{z}-i\frac{t_{x}}{2}\sigma_{x},~~T_{x}^{\dagger}\!=\!-\frac{m_{0}}{2}\sigma_{z}+i\frac{t_{x}}{2}\sigma_{x},$ (27)
$T_{y}\!=\!-\frac{m_{0}}{2}\sigma_{z}-i\frac{t_{y}}{2}\sigma_{y},~~T_{y}^{\dagger}\!=\!-\frac{m_{0}}{2}\sigma_{z}+i\frac{t_{y}}{2}\sigma_{y},$ (28)
where $z\!=\!j_{z}a$.
The detailed derivation of the real-space disordered Weyl-like model Hamiltonian in Eq. (23) with PBCs in the $x$- and $y$-directions is provided in Section SVI in the Supplemental Material Sup
IV.1 Local Chern marker
Figure 8: Local Chern marker [Eq. (29)] of the real-space disordered Weyl-like model Hamiltonian in Eq. (23), evaluated at fixed y=0y\!=\!0 and z=16az\!=\!16a (the center position along the $z$ direction with $L_{z}\!=\!\lambda_{z}\!=\!31a$), for different disorder strengths $U$. (a) $U/t_{0}\!=\!1$, (b) $U/t_{0}\!=\!2$, (c) $U/t_{0}\!=\!3$, and (d) $U/t_{0}\!=\!4$. The data with error bars represent the standard deviation of the local Chern marker calculated over 1000 samples. The parameters used are $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, $L_{x}\!=\!L_{y}\!=\!21a$, and a=1a\!=\!1 nm.
To investigate the topological invariant in disordered systems, we employ the local Chern marker, defined as Bianco and Resta (2011); Caio et al. (2019); Li et al. (2021b)
${\cal C}(m)\!=\!-\frac{4\pi}{A_{c}}{\rm Im}\sum_{s=A,B}\langle{\bf r}_{m,s}|\hat{P}\hat{x}\hat{Q}\hat{y}\hat{P}|{\bf r}_{m,s}\rangle,$ (29)
where $A_{c}$ is the area of a real-space unit cell, and $|{\bf r}_{m,s}\rangle$ denotes the sublattice state $s$ (s=A,Bs\!=\!A,B) in the $m$th unit cell, labeled by $m\!=\!(j_{x},j_{y})$. The operator $\hat{P}\!=\!\sum_{n}|\psi_{n}\rangle\langle\psi_{n}|$ is the projector onto the chosen quasienergy band in real space, while $\hat{Q}\!=\!\hat{I}\!-\!\hat{P}$ is its complement, with $\hat{I}$ being the identity operator. The operators $\hat{x}$ and $\hat{y}$ represent the position operators along the $x$ and $y$ directions, respectively.
Furthermore, we plot the local Chern marker for the real-space disordered Weyl-like model Hamiltonian in Eq. (23), evaluated at fixed y=0y\!=\!0 and z=16az\!=\!16a (the center position along the $z$ direction with $L_{z}\!=\!\lambda_{z}\!=\!31a$), for various disorder strengths $U$, as shown in Fig. 8. By comparing Figs. 8(a) and 8(b), we observe that the local Chern numbers ${\cal C}(x)$ in the bulk of the system fluctuate around a macroscopic average of ${\cal C}(x)\!=\!1$ for disordered cases with weak to moderate disorder (up to $U/t_{0}\!\approx\!2$). Particularly, the panels demonstrate that the value of the local Chern marker remains stable at the center of the $(x,y)$ plane, i.e., at (x=0,y=0)(x\!=\!0,y\!=\!0). Therefore, in the following discussions, we focus on the local Chern marker at this central point.
IV.2 Layer Hall conductance with disorder
Figure 9: Layer Hall conductances of the real-space disordered Weyl-like model Hamiltonian in Eq. (23) for different disorder strengths $U$. The two green vertical dashed lines indicate the positions of the Weyl-like points. The data with error bars represent the standard deviation of the Hall conductance calculated over 1000 samples. (a) $U/t_{0}\!=\!1$, (b) $U/t_{0}\!=\!2$, (c) $U/t_{0}\!=\!3$, and (d) $U/t_{0}\!=\!4$. Here, we use the local Chern marker evaluated at x=y=0x\!=\!y\!=\!0. The system length along the $z$ direction is $L_{z}\!=\!\lambda_{z}\!=\!31a$, and all other parameters are the same as those in Fig. 8.
Based on the local Chern marker in Eq. (29), we plot the layer Hall conductances of the real-space disordered Weyl-like model Hamiltonian in Eq. (23) for various disorder strengths, as shown in Fig. 9.
Figure 9 reveals that, although the fluctuations increase with growing disorder strength, the two Weyl-like points remain robust under weak to moderate disorder (up to $U/t_{0}\!\approx\!2$). Moreover, the Hall conductance exhibits pronounced fluctuations near the topological phase transition points, i.e., the Weyl-like points $z_{c\pm}$, even at fixed weak or moderate disorder strengths ($U/t_{0}\!\leqslant\!2$).
IV.3 3D Hall conductivity with disorder
Figure 10: 3D Hall conductivity [Eq. (8)] as a function of the disorder strength $U$ for different system sizes $L_{||}\!=\!9a$, $15a$, $21a$, and $27a$, where $L_{||}\!=\!L_{x}\!=\!L_{y}$. The data with error bars represent the standard deviation of the 3D Hall conductivity calculated over 1000 samples. The system length along the $z$-direction is $L_{z}\!=\!\lambda_{z}\!=\!61a$, and all other parameters are the same as those used in Fig. 9.
To investigate the critical disorder threshold at which the topological phase becomes unstable, we focus on the 3D Hall conductivity as a function of the disorder strength $U$, for different system sizes $L_{||}\!=\!L_{x}\!=\!L_{y}$ in the $x$ and $y$ directions, as shown in Fig. 10.
Interestingly, an inverse point is observed at $U_{\rm inv}/t_{0}\!\approx\!3.9$ (the green vertical dashed line) across different system sizes $L_{||}$, as shown in Fig. 10. For $U\!<\!U_{\rm inv}$, the 3D Hall conductivity increases with system size $L_{||}$, whereas for $U\!>\!U_{\rm inv}$, it decreases as $L_{||}$ grows. In the latter case, as $L_{||}$ becomes large, the 3D Hall conductivity approaches zero, indicating that beyond this disorder strength $U_{\rm inv}$, the topological phase tends to become unstable.
IV.4 Born approximation
Figure 11: 3D Hall conductivity [Eq. (8)] as a function of the disorder strength $U$. The black solid curve denotes the result obtained using the Born approximation [Eq. (IV.4)]. The data with error bars represent the standard deviation of the 3D Hall conductivity calculated over 1000 samples. All other parameters are the same as those used in Fig. 10.
To support the physical interpretation of the numerical simulations for the 3D Hall conductivity, as shown in Fig. 10, we analyze the present model using an effective medium theory based on the Born approximation, which neglects higher-order scattering processes Groth et al. (2009); Chen et al. (2017a, b, 2018a, 2018b); Yi et al. (2024); Chen and Zhou (2025). Within the self-consistent Born approximation, the self-energy $\Sigma$ for a finite disorder strength is determined by the following integral equation Ominato and Koshino (2014); Song et al. (2012); Hung et al. (2016); Guo et al. (2010); Groth et al. (2009); Chen et al. (2017a, b, 2018a, 2018b); Yi et al. (2024); Chen and Zhou (2025):
$\Sigma(z)\!=\!\frac{U^{2}}{3}\left(\frac{a}{2\pi}\right)^{d}$
$\times\!\!\!\int_{\rm BZ}\!\!\left\{\!\sigma_{s}\!\left[(E_{F}\!+\!i\delta)\mathbb{I}\!-\!{\cal H}({\bf k}_{||},z)\!-\!\Sigma(z)\right]^{-1}\!\sigma_{s}\!\right\}d^{2}{\bf k}_{||},$
[…]
$\Sigma(z)\!\approx\!\frac{U^{2}}{3}\left(\frac{a}{2\pi}\right)^{2}\!\!\!\int_{\rm BZ}\!\left[(E_{F}\!+\!i\delta)\mathbb{I}\!-\!{\cal H}({\bf k}_{||},z)\right]^{-1}d^{2}{\bf k}_{||}.$
The Hamiltonian matrix ${\cal H}({\bf k}_{||},z)\!=\!\sum_{i=x,y,z}h_{i}\sigma_{i}$ is $2\times 2$, so that the integrand in Eq. (IV.4) is written as Hung et al. (2016)
$\left[(E_{F}\!+\!i\delta)\mathbb{I}\!-\!{\cal H}({\bf k}_{||},z)\right]^{-1}\!=\!\frac{(E_{F}\!+\!i\delta)\mathbb{I}\!+\!{\cal H}({\bf k}_{||},z)}{(E_{F}\!+\!i\delta)^{2}\!-\!\sum_{i=x,y,z}h_{i}^{2}},$
which is more convenient for our numerical calculations.
By analyzing the renormalized Hamiltonian ${\cal H}({\bf k}_{||},z)\!+\!\Sigma(z)$, we find that the Born approximation results agree well with the numerical calculations in the weak to moderate disorder regime ($U/t_{0}\!\leqslant\!2$), as shown in Fig. 11, where the black solid line denotes the prediction from the self-consistent Born approximation.
V Floquet Hamiltonian
@@
The optical field propagating along the $z$ direction in the material is given by ${\bf E}(t)\!=\!\partial{\bf A}(t)/\partial t\!=\!E_{0}(\cos(\omega t),\cos(\omega t\!+\!\varphi),0)$, where $E_{0}$ is the amplitude, $\omega$ is the angular frequency, and $\varphi$ represents the phase. For linear polarization, $\varphi\!=\!0$, while $\varphi\!=\!\mp\pi/2$ corresponds to left- or right-handed circular polarization. The vector potential is periodic with period $T\!=\!2\pi/\omega$, and it is given by ${\bf A}(t)\!=\!{\bf A}(t\!+\!T)\!=\!\omega^{-1}E_{0}(\sin(\omega t),\sin(\omega t\!+\!\varphi),0)$. In the off-resonant regime, where the central Floquet band is well-separated from other replicas, the high-frequency expansion applies. We choose the optical frequency as $\hbar\omega\!=\!10$ eV, which is much larger than the bandwidth of the material Zhan et al. (2024); Qin et al. (2023a, 2022b, 2022a, 2024a); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018).
@@
${\cal H}^{(F)}({\bf k})\!=\!\frac{i}{T}\ln\left\{\mathcal{T}\exp\left[-i\int^{T}_{0}{\cal H}({\bf k},t)dt\right]\right\},$ (34)
where $\mathcal{T}$ is the time-ordering operator. In the high-frequency regime, a closed-form solution exists via the high-frequency expansion Magnus (1954); Blanes et al. (2009); Lee et al. (2018); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, a); Du et al. (2022); Wang et al. (2023)
@@
where ${\cal H}_{m-m^{\prime}}\!=\!\frac{1}{T}\int_{0}^{T}{\cal H}({\bf k},t)\exp[i(m-m^{\prime})\omega tdt]$, and $m$, $m^{\prime}$ are integers.
V.1 Floquet Weyl-like model
@@
where ${\cal J}_{n}(A_{0}a)$ is the $n$th Bessel function of the first kind Temme (1996), and $A_{0}\!=\!eE_{0}/(\hbar\omega)$. The detailed derivations leading to Eq. (36) are provided in Subsection SVII A in the Supplemental Material Sup . For $\varphi\!=\!0$ (linearly polarized laser), the Weyl-like nodes are determined by solving the equation $m_{z}\!+\!V_{z}\cos(Q_{z}z)\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\!=\!0$, subject to the condition $1\!\leqslant\!\cos(Q_{z}z)\!=\!\{m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\}/V_{z}\!\leqslant\!1$. This yields the Weyl-like nodes at
@@
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $1\!\leqslant\!\tilde{m}_{z}(A_{0})/V_{z}\!\leqslant\!1$. By substituting the model parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and a=1a\!=\!1 nm into the condition $|\tilde{m}_{z}(A_{0})/V_{z}|\!<\!1$, we have $A_{0}\!<\!1.034$ nm-1. From Eq. (40), it is clear that the positions of the Weyl-like nodes can be tuned by adjusting the intensity of the optical field.
Furthermore, substituting Eq. (40) into Eq. (9), we can obtain the total 3D Hall conductivity for the Floquet Weyl-like model in Eq. (36) as
@@
Additionally, Section SVIII in the Supplemental Material Sup presents the derivations for the Floquet Weyl-like model under PBCs along the $x$ direction and OBCs along the $y$ direction.
V.2 Floquet Weyl semimetal
The Floquet Hamiltonian for the Weyl semimetal, derived from Eq. (35), takes the form
@@ V.2 Floquet Weyl semimetal
A detailed derivation of Eq. (42) is provided in Subsection SVII B in the Supplemental Material Sup . For linearly polarized light ($\varphi\!=\!0$), the Weyl nodes are determined by solving the equation $m_{z}\!+\!t_{z}\cos(k_{z}a)\!-\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\!=\!0$, subject to the condition $1\!\leqslant\!\cos(k_{z}a)\!=\!\{m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]\}/t_{z}\!\leqslant\!1$. This condition gives the Weyl node positions
@@
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $1\!\leqslant\!\tilde{m}_{z}(A_{0})/t_{z}\!\leqslant\!1$. Substituting typical model parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and a=1a\!=\!1 nm into the condition $|\tilde{m}_{z}(A_{0})/t_{z}|\!<\!1$, we obtain the constraint $A_{0}\!<\!1.034$ nm-1. From Eq. (46), we can conclude that the positions of the Weyl nodes can be tuned by adjusting the intensity of the optical field.
Furthermore, substituting Eq. (46) into Eq. (15), we can obtain the total 3D Hall conductivity for the Floquet Weyl semimetal in Eq. (42) as
@@
The detailed derivation of the Floquet Hamiltonian matrix for the Weyl semimetal, under OBCs along the $z$ direction and PBCs along the $x$ and $y$ directions, is provided in Subsection SIX A in the Supplemental Material Sup . Similarly, Subsection SIX B in the Supplemental Material Sup presents the derivation for the case of PBCs along the $x$ and $z$ directions and OBCs along the $y$ direction.
V.3 Validity